# Is It Really You? Exploring Biometric Verification Scenarios in Photorealistic Talking-Head Avatar Videos

### v1

@@ Abstract
Photorealistic talking-head avatars are becoming increasingly common in virtual meetings, gaming, and social platforms. These avatars allow for more immersive communication, but they also introduce serious security risks. One emerging threat is impersonation: an attacker can steal a user’s avatar—preserving their appearance and voice—making it nearly impossible to detect its fraudulent usage by sight or sound alone. In this paper, we explore the challenge of biometric verification in such avatar-mediated scenarios. Our main question is whether an individual’s facial motion patterns can serve as reliable behavioral biometrics to verify their identity when the avatar’s visual appearance is a facsimile of its owner. To answer this question, we introduce a new dataset of realistic avatar videos created using a state-of-the-art one-shot avatar generation model, GAGAvatar, with genuine and impostor avatar videos. We also propose a lightweight, explainable spatio-temporal Graph Convolutional Network architecture with temporal attention pooling, that uses only facial landmarks to model dynamic facial gestures. Experimental results demonstrate that facial motion cues enable meaningful identity verification with AUC values approaching 80%. The proposed benchmark and biometric system are available [^1] https://github.com/BiDAlab/GAGAvatar-Benchmark for the research community in order to bring attention to the urgent need for more advanced behavioral biometric defenses in avatar-based communication systems.
@@ 1 Introduction
In recent years, there has been a surge in novel methods for generating photorealistic avatars [3, 12, 13, 21] that can create and animate high-quality 3D human avatars with a single image or text prompt. In parallel, industry investment in avatar technology has exploded, with companies such as Synthesia [^2] https://cnb.cx/3WfCHFt securing hundreds of millions of dollars in funding to scale avatar services which are being used by over one million users. Another noteworthy example is Meta [^3] https://www.uploadvr.com/meta-project-warhol-codec-avatars-training-paying/ offering to pay users to capture their facial expressions, speech and gestures to train its Codec Avatars model [15].
These photorealistic talking-head avatars are becoming popular in virtual meetings, gaming and metaverse platforms. A good example of this popularity was the virtual-reality interview [^4] https://www.youtube.com/watch?v=MVYrJJNdrEg between Mark Zuckerberg and Lex Fridman in his podcast, which had nearly 3 million views. While these avatars allow for immersive communication, they also pose critical security risks for society, as depicted in Fig. 1. An impostor can steal the avatar of an identity and impersonate him or her in real time [18].
@@
A recent example of such identity theft was a video scam in early 2024555https://edition.cnn.com/2024/02/04/asia/deepfake-cfo-scam-hong-kong-intl-hnk, in which criminals used an AI-generated video of a company executive to fool another employee into transferring $\approx$\$25M USD to fraudulent accounts. This incident demonstrates just how difficult it is to distinguish a real person from an impersonated one. Indeed, as avatar technology improves, such impostor attacks may become nearly impossible to detect by sight or sound alone. This is the specific use case considered in this study, where the avatar stolen by the impostor is so realistic in terms of appearance and voice that only behavioral biometric information, such as facial gestures, can be used to detect if the driver identity moving the avatar is the real owner of the avatar or an impostor.
It is important to highlight that the aforementioned scenario is different than the traditional DeepFake detection scenario [17, 19, 20], in which the main task consists of simply predicting if a video is real or fake, not whether the person driving the avatar is, in fact, the owner of the avatar.
@@
Can an individual’s facial motion patterns serve as a reliable biometric characteristic to verify their identity during avatar-mediated communication?
In the context of facial avatars, each person’s manner of speaking and emoting is distinctive and can be difficult to imitate. By capturing these subtle dynamics, it might be possible to verify if a photorealistic talking-head avatar is being driven by its owner or an impostor.
To investigate this question, we design an end-to-end experimental framework that simulates real-world avatar-mediated impersonation attacks while isolating facial motion as the distinguishing biometric signal. To the best of our knowledge, the only study that has previously analyzed this scenario is [16], from NVIDIA. In their study, the authors coined the term “avatar fingerprinting” to refer to the detection of unauthorized use of avatars. While their study reported promising results, we encountered the following problems when attempting to reproduce it: i) the proposed method and its corresponding code is not publicly available, limiting its reproducibility, and ii) most research groups do not have access to the GPU resources required to train NVIDIA’s verification model. Our study aims to bridge these gaps by publicly releasing our proposed avatar dataset and standard benchmark, as well as exploring lighter biometric verification approaches.
@@
  • A new public avatar video dataset generated using a state-of-the-art one-shot avatar generator, GAGAvatar [3], including both genuine and impostor avatar videos.
@@
  • A first exploration of this challenging avatar scenario proposing a lightweight, explainable behavioral biometric system that is based solely on the facial motion patterns of the identities. Our approach is based on a Graph Convolutional Network (GCN). Unlike [16] that relies on high-dimensional pairwise distance tensors (i.e., all-pairs landmark distances), our approach is well suited for facial motion modeling as it explicitly encodes the mesh-like geometry of the face through graph structure. This results in both spatial awareness (by preserving local neighborhood relationships among facial regions) and parameter efficiency, given the small and fixed graph size typical of landmark-based representations, allowing faster training and inference.
@@ 3.1 Real datasets
We use two public video datasets for avatar video generation. These datasets are selected as they simulate a similar environment to virtual meetings.
  • CREMA-D [^6] https://github.com/CheyneyComputerScience/CREMA-D [2]: This dataset provides 7,442 short videos of 91 actors (48 male, 43 female, ages 20–74, diverse ethnicity) performing scripted sentences in various emotional states. Each actor utters 12 different sentences with one of six basic emotions (anger, disgust, fear, happy, neutral and sad) at multiple intensity levels. These videos were recorded in a controlled setting with frontal views of the speaker.
  • RAVDESS [^7] https://zenodo.org/records/1188976 [14]: This dataset contains 24 actors (12 female, 12 male) speaking two fixed statements with eight emotions (calm, happy, sad, angry, fearful, surprise, disgust and neutral). All videos were recorded in a studio environment with actors facing front against a green-screen background. In total, we use the 1,440 video-only speech files from the original dataset.
@@ 3.2 Avatar Video Dataset Generation
To generate our avatar dataset, we employ the state-of-the-art avatar generation model GAGAvatar [^8] https://github.com/xg-chu/GAGAvatar [3]. This model takes a reference image (representing the target identity’s appearance) and a driving video (providing the driver identity’s motion), and generates a new avatar video that preserves the visual appearance of the reference image while reproducing the facial expressions and movements from the driving video. The resolution of the videos generated is 512x512 with a black background. It is worth noting that the GAGAvatar model does not use the audio from the original videos to generate the avatars, only the visual data.
@@
Despite GAGAvatar being state-of-the-art, in some cases it cannot generate avatar videos that perfectly match the expressions from the driving video. For instance, in Fig. 2 (left), the column corresponding to “Driver $i$ shows that, for the avatar generated with appearance from “Target $j$, the facial expression is subtly different from the original video. The avatar is not frowning as much, and its mouth does not follow the original expression with precision.
@@ 4.1 Landmarks Extraction
We then normalize the facial landmarks using a frame-based approach. In particular, we take the 3D position of the nose tip landmark shown in red in Fig. 4 (c), and we subtract it from all the other landmark positions in that frame to achieve translation invariance. Then, we scale the landmarks positions with the intercanthal distance, i.e., distance between the two inner corners of the eyes, shown in blue in Fig. 4 (c), resulting in scale invariance.
@@ 4.2 Graph Neural Networks
A Graph Neural Network (GNN) is a neural architecture designed to process data represented as a graph $G=(V,E)$, where $V$ is the set of nodes and $E$ the set of edges. Each node $v\in V$ is associated with an initial feature vector $x_{v}\in\mathbb{R}^{d}$ (i.e., a landmark’s 3D coordinate, with $d=3$) and, through a series of $L$ message–passing layers, the GNN updates each node’s features $h_{v}^{(l)}$ by aggregating information from its neighbors, i.e., nodes connected to it by an edge.
By stacking $L$ layers, each node’s representation $h_{v}^{(L)}$ incorporates information from nodes up to $L$ hops away, capturing both local and global graph structure. The final node embeddings can then be pooled (i.e., averaged or summed) to produce a graph‐level embedding, $h_{G}$.
GNNs are particularly well suited for facial motion modeling because they explicitly encode the mesh-like geometry of the face through graph structure, enabling parameter sharing across spatially connected landmarks. This results in both spatial awareness (by preserving local neighborhood relationships among facial regions) and parameter efficiency, given the small and fixed graph size typical of landmark-based representations. In contrast to approaches that rely on high-dimensional pairwise distance tensors (e.g., all-pairs landmark distances as in [16]), a graph representation avoids quadratic growth in input size and facilitates faster training and inference.
@@
These unique properties make GCNs an effective choice for learning per-frame representations that are sensitive to the spatio-structural patterns characteristic of an individual’s facial motion.
@@ 4.3 Proposed System
Our biometric system follows a spatio-temporal architecture designed to learn identity‐specific facial motion patterns from avatar videos. As illustrated in Fig. 3, the core idea is to represent each video as a sequence of facial landmark graphs, process these with a GCN to encode spatial structure, and then aggregate them via a temporal attention-based pooling mechanism.
Each avatar video is divided into non-overlapping fixed-length clips of $T=50$ frames. This choice is motivated by prior work [16], which found that sequences approaching $\approx$30 frames begin to yield strong verification performance with landmark-based inputs.
Given an avatar clip of length $T$ frames, we first extract $V=109$ 3D facial landmarks per frame $t$ using MediaPipe [11]. For each frame $t$, the landmarks define the nodes $V$ of a graph whose edges $E$ are constructed via Delaunay triangulation [6]. The result is a sequence $\{G^{(t)}\}_{t=1}^{T}$ of graphs that encode both local and global facial structure.
Each graph $G^{(t)}$ in the sequence is then processed by a 3-layer GCN ($L=3$), in which the first layer transforms the initial feature vector $x_{v}$ into an embedding $h_{v}^{(1)}\in\mathbb{R}^{d^{(1)}}$, the second layer transforms $h_{v}^{(1)}$ into $h_{v}^{(2)}\in\mathbb{R}^{d^{(2)}}$, and the third layer transforms $h_{v}^{(2)}$ into $h_{v}^{(3)}\in\mathbb{R}^{d^{(3)}}$, being $d^{(1)}=d^{(2)}=64$, and $d^{(3)}=256$. Within each layer, node features are updated by aggregating information from their immediate neighbors, followed by non-linear ReLU activation and dropout ($p=0.3$). Applying the same GCN across all $T$ frames results in a sequence of spatial embeddings $\{h_{G}^{(t)}\}_{t=1}^{T}$ that encode the per-frame geometry of facial expressions.
To convert this sequence of frame embeddings into a single, discriminative descriptor of the entire clip, $e\in\mathbb{R}^{d^{(3)}}$, we consider a temporal attention mechanism. This module learns to assign weights to each frame-level embedding based on its relevance for identity verification, yielding higher attention weights in frames with distinctive facial motion patterns and smaller attention weights in less distinctive intervals (e.g., static segments). The attention scores are applied to the frame embeddings to compute a weighted sum, yielding the final clip‐level embedding.
@@ 5 Experimental Protocol
For reproducibility reasons, we adopt the same train-validation-test split on CREMA-D and RAVDESS datasets as in [16]. In this protocol, all identities are split disjointly across sets to ensure that no identity in the test set (neither targets nor drivers) are seen during training or validation. The same identity can only be used as driver or target within the same split. We refer to these data splits as $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{val}}$, and $\mathcal{D}_{\text{test}}$. Next, we describe the key details for the training and final evaluation protocols.
@@ 5.1 Biometric System: Training
Our biometric system is trained on $\mathcal{D}_{\text{train}}$, with hyperparameters tuned on $\mathcal{D}_{\text{val}}$. The final model checkpoint is selected based on the lowest validation loss. Training is conducted using triplet loss. Each triplet $(a,p,n)$ consists of: i) an anchor, $a$, which can be any avatar video (genuine avatar or impostor avatar), ii) a positive sample, $p$, which is any avatar video driven by the same person as $a$ (regardless of target identity), and iii) a negative sample, $n$, which is any avatar video driven by a different person than $a$ (again, regardless of target identity). We do not mine hard or semi-hard triplets, but only random triplets.
@@ 5.2 Biometric System: Evaluation
For a given identity $i$ in $\mathcal{D}_{\text{test}}$, we use a high-quality neutral image (target image), from which we generate a genuine avatar. We use all original videos from identity $i$ to generate $N$ genuine avatar video clips, which simulate the real usage of $i$’s avatar. From all those genuine avatar clips, we take one as reference clip, and we build all the $N-1$ genuine comparisons with the remaining genuine avatar clips. We gather all other identities in $\mathcal{D}_{\text{test}}$, and we generate $M$ impostor avatar clips by using the impostor videos and the target’s avatar. We then use the same reference genuine avatar clip to compare with the $M$ impostor avatar clips, obtaining $M$ impostor comparisons.

### v2

@@ Abstract
Photorealistic talking-head avatars are becoming increasingly common in virtual meetings, gaming, and social platforms. These avatars allow for more immersive communication, but they also introduce serious security risks. One emerging threat is impersonation: an attacker can steal a user’s avatar, preserving his appearance and voice,making it nearly impossible to detect its fraudulent usage by sight or sound alone. In this paper, we explore the challenge of biometric verification in such avatar-mediated scenarios. Our main question is whether an individual’s facial motion patterns can serve as reliable behavioral biometrics to verify their identity when the avatar’s visual appearance is a facsimile of its owner. To answer this question, we introduce a new dataset of realistic avatar videos created using a state-of-the-art one-shot avatar generation model, GAGAvatar, with genuine and impostor avatar videos. We also propose a lightweight, explainable spatio-temporal Graph Convolutional Network architecture with temporal attention pooling, that uses only facial landmarks to model dynamic facial gestures. Experimental results demonstrate that facial motion cues enable meaningful identity verification with AUC values approaching 80%. The proposed benchmark and biometric system are available [^1] https://github.com/BiDAlab/GAGAvatar-Benchmark for the research community in order to bring attention to the urgent need for more advanced behavioral biometric defenses in avatar-based communication systems.
@@ 1 Introduction
In recent years, there has been a surge in novel methods for generating photorealistic avatars [5, 21, 22, 34] that can create and animate high-quality 3D human avatars with a single image or text prompt. In parallel, industry investment in avatar technology has exploded, with companies such as Synthesia [^2] https://cnb.cx/3WfCHFt securing hundreds of millions of dollars in funding to scale avatar services, which are being used by more than one million users. Another notable example is Meta [^3] https://www.uploadvr.com/meta-project-warhol-codec-avatars-training-paying/ that pays users to capture their facial expressions, speech, and gestures to train its Codec Avatars model [25].
These photorealistic talking-head avatars are becoming popular in virtual meetings, gaming, and metaverse platforms. A good example of this popularity was the virtual reality interview [^4] https://www.youtube.com/watch?v=MVYrJJNdrEg between Mark Zuckerberg and Lex Fridman in his podcast, which had nearly 3 million views. Although these avatars allow for immersive communication, they also pose critical security risks to society, as depicted in Fig. 1. An impostor can steal the avatar of an identity and impersonate him or her in real time [30].
@@
A recent example of such identity theft was a video scam in early 2024555https://edition.cnn.com/2024/02/04/asia/deepfake-cfo-scam-hong-kong-intl-hnk, in which criminals used an AI-generated video of a company executive to fool another employee into transferring $\approx$\$25M USD to fraudulent accounts. This incident demonstrates how difficult it is to distinguish a real person from an impersonated one. Indeed, as avatar technology improves, such impostor attacks may become nearly impossible to detect by sight or sound alone. This is the specific use case considered in this study, where the avatar stolen by the impostor is so realistic in terms of appearance and voice that only behavioral biometric information, such as facial gestures, can be used to detect whether the driver identity that moves the avatar is the real owner of the avatar or an impostor.
It is important to note that the above scenario is different from DeepFake detection [31, 32, 33], in which the main task consists in simply predicting if a video is real or fake, not whether the person driving the avatar is, in fact, the owner of the avatar. The problem is also different (but related) to presentation attack detection (PAD) in biometrics [17, 18]. Considering the differences, we can exploit methods and consolidated knowledge from both fields (DeepFakes and biometric PAD).
@@
Can facial motion patterns serve as a reliable biometric characteristic to verify identity during avatar-mediated communication?
In the context of avatars, each person’s manner of speaking and emoting is distinctive and can be difficult to imitate. By capturing these subtle dynamics, it might be possible to verify if a photorealistic talking-head avatar is being driven by its owner or an impostor. Capturing these subtle dynamics is based on our previous work on behavioral face analysis for e-health [11, 14] and e-learning [13, 6].
To investigate this question, we designed an end-to-end experimental framework that simulates real-world avatar-mediated impersonation attacks while isolating facial motion as the distinguishing biometric signal. To our knowledge, the only study that has previously analyzed this scenario is [29], from NVIDIA. In their study, the authors coined the term “avatar fingerprinting” to refer to the detection of unauthorized use of avatars. Although their study reported promising results, we encountered the following problems when trying to reproduce it: i) the proposed method and its corresponding code are not publicly available, limiting its reproducibility, and ii) most research groups do not have access to the GPU resources required to train NVIDIA’s verification model. Our study aims to bridge these gaps by publicly releasing our proposed avatar dataset and standard benchmark, as well as exploring lighter biometric verification approaches.
@@
  • A new public avatar video dataset generated using a state-of-the-art one-shot avatar generator, GAGAvatar [5], including genuine and impostor avatar videos.
@@
  • A first exploration of this challenging avatar scenario proposing a lightweight and explainable behavioral biometric system that is based solely on the facial motion patterns of the identities. Our approach is based on a Graph Convolutional Network (GCN). Unlike [29] which relies on high-dimensional pairwise distance tensors (i.e., all-pair landmark distances), our approach is well suited for facial motion modeling as it explicitly encodes the mesh-like geometry of the face through graph structure. This results in both spatial awareness (by preserving local neighborhood relationships among facial regions) and parameter efficiency, given the small and fixed graph size typical of landmark-based representations, allowing faster training and inference.
@@ 3.1 Real datasets
We use two public video datasets for avatar video generation. These datasets are selected because they simulate a similar environment to virtual meetings.
  • CREMA-D [^6] https://github.com/CheyneyComputerScience/CREMA-D [4]: This dataset provides 7,442 short videos of 91 actors (48 male, 43 female, ages 20–74, diverse ethnicity) performing scripted sentences in various emotional states. Each actor utters 12 different sentences with one of six basic emotions (anger, disgust, fear, happy, neutral, and sad) at multiple intensity levels [27]. These videos were recorded in a controlled setting with frontal views of the speaker.
  • RAVDESS [^7] https://zenodo.org/records/1188976 [23]: This dataset contains 24 actors (12 female, 12 male) speaking two fixed statements with eight emotions (calm, happy, sad, angry, fearful, surprise, disgust, and neutral). All videos were recorded in a studio environment with actors facing up against a green screen background. In total, we used 1,440 video-only speech files from the original dataset.
@@ 3.2 Avatar Video Dataset Generation
To generate our avatar dataset, we employ the state-of-the-art avatar generation model GAGAvatar [^8] https://github.com/xg-chu/GAGAvatar [5]. This model takes a reference image (representing the target identity’s appearance) and a driving video (providing the driver identity’s motion), and generates a new avatar video that preserves the visual appearance of the reference image while reproducing the facial expressions and movements from the driving video. The resolution of the videos generated is 512$\times$512 on a black background. It is worth noting that the GAGAvatar model does not use the audio from the original videos to generate the avatars, only the visual data.
@@
Despite GAGAvatar being state-of-the-art, in some cases it cannot generate avatar videos that perfectly match the expressions from the driving video. For example, in Fig. 2 (left), the column corresponding to “Driver $i$ shows that, for the avatar generated with the appearance of “Target $j$, the facial expression is subtly different from the original video. The avatar is not frowning as much and its mouth does not follow the original expression with precision.
@@ 4.1 Landmarks Extraction
We then normalize the facial landmarks using a frame-based approach. In particular, we take the 3D position of the nose tip landmark shown in red in Fig. 4 (c), and subtract it from all the other landmark positions in that frame to achieve translation invariance. Then we scale the landmark positions with the intercanthal distance, i.e., the distance between the two inner corners of the eyes, shown in blue in Fig. 4 (c), resulting in scale invariance.
@@ 4.2 Graph Neural Networks
A Graph Neural Network (GNN) is a neural architecture designed to process data represented as a graph $G=(V,E)$, where $V$ is the set of nodes and $E$ the set of edges [24]. Each node $v\in V$ is associated with an initial feature vector $x_{v}\in\mathbb{R}^{d}$ (i.e., the 3D coordinate of a landmark, with $d=3$) and, through a series of messages passing layers $L$, the GNN updates the features of each node $h_{v}^{(l)}$ aggregating information from its neighbors, i.e., nodes connected to it by an edge.
By stacking $L$ layers, the representation of each node $h_{v}^{(L)}$ incorporates information from nodes up to $L$ hops away, captures both the local and global graph structure. The final node embeddings can then be pooled (i.e., averaged or summed) to produce a graph‐level embedding, $h_{G}$.
GNNs are particularly well suited for facial motion modeling because they explicitly encode the mesh-like geometry of the face through a graph structure, enabling parameter sharing across spatially connected landmarks. This results in both spatial awareness (by preserving local neighborhood relationships among facial regions) and parameter efficiency, given the small and fixed graph size typical of landmark-based representations. In contrast to approaches that rely on high-dimensional pairwise distance tensors (e.g., all-pairs landmark distances as in [29]), a graph representation avoids quadratic growth in input size and facilitates faster training and inference.
@@ 4.3 Proposed System
Our biometric system follows a spatio-temporal architecture designed to learn identity‐specific facial motion patterns from avatar videos. As illustrated in Fig. 3, the core idea is to represent each video as a sequence of facial landmark graphs, process them with a GCN to encode spatial structure, and then aggregate them via a temporal attention-based pooling mechanism.
Each avatar video is divided into non-overlapping fixed-length clips of $T=50$ frames. This choice is motivated by previous work [29], which found that sequences approaching $\approx$30 frames begin to produce a strong verification performance with landmark-based input.
Given an avatar clip of length $T$ frames, we first extract $V=109$ 3D facial landmarks per frame $t$ using MediaPipe [20]. For each frame $t$, the landmarks define the nodes $V$ of a graph whose edges $E$ are constructed by Delaunay triangulation [9]. The result is a sequence $\{G^{(t)}\}_{t=1}^{T}$ of graphs that encode both local and global facial structure.
Each graph $G^{(t)}$ in the sequence is then processed by a 3-layer GCN ($L=3$), in which the first layer transforms the initial feature vector $x_{v}$ into an embedding $h_{v}^{(1)}\in\mathbb{R}^{d^{(1)}}$, the second layer transforms $h_{v}^{(1)}$ into $h_{v}^{(2)}\in\mathbb{R}^{d^{(2)}}$, and the third layer transforms $h_{v}^{(2)}$ into $h_{v}^{(3)}\in\mathbb{R}^{d^{(3)}}$, being $d^{(1)}=d^{(2)}=64$, and $d^{(3)}=256$. Within each layer, node features are updated by aggregating information from their immediate neighbors, followed by non-linear ReLU activation and dropout ($p=0.3$). Applying the same GCN across all $T$ frames results in a sequence of spatial embeddings $\{h_{G}^{(t)}\}_{t=1}^{T}$ that encode the geometry of facial expressions.
To convert this sequence of frame embeddings into a single discriminative descriptor of the entire clip, $e\in\mathbb{R}^{d^{(3)}}$, we consider a temporal attention mechanism. This module learns to assign weights to each embedding at the frame level based on its relevance to identity verification, resulting in higher attention weights in frames with distinctive facial motion patterns and smaller attention weights in less distinctive intervals (e.g., static segments). Attention scores are applied to frame embeddings to compute a weighted sum, resulting in the final clip-level embedding.
@@ 5 Experimental Protocol
For reproducibility reasons, we adopt the same train-validation-test split on the CREMA-D and RAVDESS datasets as in [29]. In this protocol, all identities are split disjointly between sets to ensure that no identity is seen in the test set (no targets or drivers) during training or validation. The same identity can only be used as a driver or target within the same split. We refer to these data splits as $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{val}}$, and $\mathcal{D}_{\text{test}}$. Next, we describe the key details for the training and final evaluation protocols.
@@ 5.1 Biometric System: Training
Our biometric system is trained on $\mathcal{D}_{\text{train}}$, with hyperparameters tuned on $\mathcal{D}_{\text{val}}$. The final model checkpoint is selected on the basis of the lowest validation loss. Training is carried out using triplet loss [26]. Each triplet $(a,p,n)$ consists of: i) an anchor, $a$, which can be any avatar video (genuine avatar or impostor avatar), ii) a positive sample, $p$, which is any avatar video driven by the same person as $a$ (regardless of target identity), and iii) a negative sample, $n$, which is any avatar video driven by a different person than $a$ (again, regardless of target identity). We do not mine hard or semi-hard triplets, but only random triplets.
@@ 5.2 Biometric System: Evaluation
For a given identity $i$ in $\mathcal{D}_{\text{test}}$, we use a neutral high quality image (target image), from which we generate a original avatar. We use all original videos from identity $i$ to generate $N$ genuine avatar video clips, which simulate the real usage of avatar $i$. From all those genuine avatar clips, we take one as reference clip, and we build all the $N-1$ genuine comparisons with the remaining genuine avatar clips. We gather all other identities in $\mathcal{D}_{\text{test}}$, and we generate $M$ impostor avatar clips using the impostor videos and the target avatar. We then use the same reference genuine avatar clip to compare with the $M$ impostor avatar clips, obtaining $M$ impostor comparisons.